percentage for math last term 85% if my score out of 20 marks for the first nine tests were 14 18 16 16 18 15 20 and 19 find my score for the last test out of 20
Answer by JulietG(1812)

Total points available are 10 * 20 = 200
Note that you only have 8 scores noted. Let's call the 9th a 16 (change the equation value if it's more/less)
Total points already scored = 14+18+16+16+18+15+20+19+16= 152
Points you want to have = .85 * 200 = 170
Points you need to get on next test = 18 (170-152)